Mary's age is 3 times as old as her son. In 12 years Mary's age will be one year less than twice her son's age.How old is each now?
Answer by ewatrrr(24785)


Hi
Mary's age is 3 times as old as her son
Let x and 3x represent the son's and Mary's age NOW respectively
Question states*** In 12 years...
3x + 12 = 2(x+12) -1
Solving for x
x = 24-1-12
x = 11, son's age now. Mary is 33
CHECKING our Answer***
45 = 2*23 - 1